# Universal Vessel Segmentation for Multi-Modality Retinal Images

### v1

@@ Abstract
We identify two major limitations in the existing studies on retinal vessel segmentation: (1) Most existing works are restricted to one modality, i.e, the Color Fundus (CF). However, multi-modality retinal images are used every day in the study of retina and retinal diseases, and the study of vessel segmentation on the other modalities is scarce; (2) Even though a small amount of works extended their experiments to limited new modalities such as the Multi-Color Scanning Laser Ophthalmoscopy (MC), these works still require finetuning a separate model for the new modality. And the finetuning will require extra training data, which is difficult to acquire. In this work, we present a foundational universal vessel segmentation model (UVSM) for multi-modality retinal images. Not only do we perform the study on a much wider range of modalities, but also we propose a universal model to segment the vessels in all these commonly-used modalities. Despite being much more versatile comparing with existing methods, our universal model still demonstrates comparable performance with the state-of-the-art finetuned methods. To the best of our knowledge, this is the first work that achieves cross-modality retinal vessel segmentation and also the first work to study retinal vessel segmentation in some novel modalities.  [^1] Our codes and model will be available at the time of publication. We are also actively looking for approaches to release the three new vessel segmentation datasets. However, since retinal images and retinal vessels contain sensitive personal information, we can not make promise at this moment.
@@
retina, retinal images, vessel segmentation, image translation, topology-preserving image segmentation.
@@ I Background and Introduction
Retinal vessel segmentation is a fundamental task in retinal image processing. It has multiple useful applications, including retinal image registration [1, 2], human biometric identification [3], retinal artery-vein classification [4] and tree topology analysis in digital images [5]. Clinically, the evaluation of retinal vessels is also important since the vessel features are predictive in several systemic diseases [6, 7, 8] and vessel segmentation is the key towards the automation of the evaluation. However, most works in retinal vessel segmentation focus on only the Color Fundus (CF) images. CF uses white visible light to illuminate the retina and the reflected light is used for generating the image. Nevertheless, although CF was the most widely-used modality in the past, it requires dilation before the imaging and is uncomfortable for the patients. Additionally, the pathological information provided by CF is limited. Therefore, in more recent years, CF is less and less used and emerging new imaging techniques become increasingly popular. These modalities include (Fig. 1): (1) Multi-Color Scanning Laser Ophthalmoscopy (MC), which uses more intense laser beam and allow exposing structures in deeper layers of the retina; (2) Florescence Angiography (FA), which uses florescence dye to illuminate the retinal vessels and are commonly used in the diagnosis of diseases related to neovascularization; (3) Fundus AutoFlorescence (FAF), which does not rely on external light source but use natural fluorescence of the retina and provide insights into the health and metabolism of the retinal pigment epithelium (RPE) and photoreceptor cells; (4) Infrared/Near-Infrared Reflectance (IR/NIR), which uses Infrared or Near-Infrared lights to illuminate the image and allows seeing deeper features on the retina such as RPE abnormalities. (5) Optical Coherence Tomograph Angiograph (OCTA), which uses low-coherence interferometry to measure the backscattered light from retinal layers. It is primarily used to image the micro-vascular structure on the retina. Nevertheless, OCTA has a much smaller imaging area than the other modalities and has significantly different retinal vessel morphology and topology. Therefore, in this work we focus on the first four modalities and the CF.
@@
Figure 1: Overview of multi-modality retinal images and the results from our UVSM. In each image, the upper half is the original image and the lower half is the segmentation. In the middle, the camera (left) and modality (right) of the image is labeled. ‘Heidel’ refers to Heidelberg Spectralis.
@@
One major challenge that impedes the development of machine learning-based retinal vessel segmentation algorithms is the lack of data. Due to the complexity of the retinal vessels and the existence of many fine details, labeling is very time-consuming. Existing retinal vessel segmentation datasets (especially considering high label quality) are scarce and small even for the CF modality, which usually only have 20-40 images. In addition, to the best of our knowledge, there are no publicly available vessel segmentation datasets for FAF and IR. For FA, there is only a very small public dataset with 8 images [9]. In this work, we prepare three high-quality datasets for FA, FAF and IR, respectively. Each dataset has 40 images and the vessels are professionally and carefully labeled. Nevertheless, despite the availability of the new datasets, in this work, we only use them for evaluation purpose. The reason is that: we believe that there can never be enough annotated images to allow training a powerful universal model. Instead, we need to train the model with limited existing data and seek to overcome the challenge through innovations in the algorithm. In fact, our segmentation network is trained with only 19 annotated CF images from the widely-used DRIVE [10] dataset.
@@
There are limited previous works which are extended beyond CF modality [11, 12, 13]. However, not only they did not work with as comprehensive modalities as we do, but also in these work one still needs to finetune a separate model on the target modality and each model only works for one modality. We instead aim to build a universal model to use a single model and allows robust segmentation for all the commonly-used modalities. The model will serve as an important tool in cross-modality retinal image study, which becomes increasingly important in modern research of the retina. To achieve the universal segmentation with limited labeled data. We propose our model in two major steps (Fig. 2). Firstly, we propose to use image translation as a preliminary domain adaptation method and train a model that translate input images from arbitrary modalities and cameras into a uniform Topcon (The Topcon TRC-50DX series is explicitly referred to as the gold standard in color fundus photography) CF image modality. We accessed abundant of de-identified multi-modality retinal images from multiple widely-used cameras (Topcon, Heidelberg Spectralis, Optos) at the Jacobs Retina Center at Shiley Eye Institute, University of California, San Diego (UCSD). Those images allow us to train a useful image translation model in a self-supervised approach. Furthermore, we adopt data augmentation to account for the slight domain difference with the other CF cameras which our center does not have, including Canon and Zeiss. Secondly, we train a segmentation model on the Topcon CF domain to perform the segmentation. However, image translation can not fully eliminate the domain gaps between different modalities given such a difficult many-to-one translation task. To solve this problem, we propose to use a topology-aware segmentation method that learns topological features of the vessels in addition to their conventional pixel-level features. The topology-aware segmentation methods were initially proposed to improve the topological accuracy of the segmentation of tubular structures within a certain type of image. We extend these methods to our task and found that not only do they improve the topological accuracy in the segmentation (e.g., preserving vessel continuity, branching and loop features), but also they force the segmentation to be topologically correct so when images with slight domain gap are inputted, the model can still make the correct segmentation.
@@
Our universal vessel segmentation model is evaluated on 7 diverse datasets (including 3 new datasets we prepared for this work) of the 5 most commonly-used retinal image modalities (CF,FA,FAF,MC,IR/NIR) taken from 5 most commonly-used cameras (Topcon, Canon, Zeiss, Heidelberg Spectralis, Optos). Our universal model is compared with state-of-the-art methods which are finetuned on each of the 7 datasets, respectively, and is shown to achieve comparable performance. Extensive ablation studies were also performed to investigate how different design choices in the proposed pipeline affects the performance of the universal vessel segmentation. The main contributions of this paper can be summerized as:
@@ II-A Retinal Vessel Segmentation
Retinal vessel segmentation classifies each pixel on the retinal images into foreground (vessel) and background (non-vessel). And the literature focuses on segmenting vessels in the CF images. Various machine learning-based or non-machine learning-based methods were proposed. A majority of the early methods are based on the matched filter [14, 15, 16], which convolves retinal images with predefined or learned kernels and then adaptively threshold the image to obtain the vessel segmentation map. Another type of methods use tracking algorithms to mathematically estimate the growth of the retinal vessel trees from seed points [17, 18]. The segmentation can then be derived by extending around the estimated vessel centerlines using the estimated vessel diameters. Some methods also adopt morphological image processing methods such as the top-hat operation to extract the vessel or as a post-processing step [19, 20]. Furthermore, numerous traditional machine learning-based methods were also applied in retinal vessel segmentation. Many classic models were used, such as the support vector machine [21], random forest [22] and Adaboost [23] for supervised methods, and Gaussian mixture model [24], Fuzzy C-Means [25] for unsupervised methods. In more recent years, deep learning-based methods dominate the literature. They majorly rely on the Convolutional Neural Networks (CNN) as a backbone model due to the existence of vast fine details and the whole-spreading characteristic of the vessels. An introductory work is the Deep Retinal Image Understanding (DRIU) [26], which adopts a VGG [27] network structure as encoder and assign specialized output layers to the embedded features to segment the vessels and the optic disc from the retina. Based on a widely-used U-Net [28] network structure and to seek further improvement for retinal vessel segmentation, [29, 30, 31, 32, 33] proposed to add various global reasoning modules at the bottom of the network. [32, 33, 34, 35, 31] proposed to add feature aggregation modules in the residual connection at each level of the network. [36, 37] proposed to use deformable convolution where the convolution kernel shapes are learned to adapt to the elongated morphology of the retinal vessels. With similar motivation, [38] proposed an orientation-selective convolution to learn selective receptive fields of the kernel to adapt to the shape of the vessels. [31] proposed to use improved feature fusion methods to better preserve the fine details of vessels in the segmentation. In addition to a single U-Net-like segmentation network, several works [39, 38, 40] proposed to use cascaded networks or a single looped network to iteratively refine the segmentation.
@@ II-C Topology-Aware Image Segmentation
Topology-aware image segmentation aims to preserve the topological accuracy in the segmentation of tubular and net-like structures. The literature mainly researches a topological loss function that extracts topological features from the ground truth and segmentation and minimizes the difference between them. One type of methods extract certain features from the segmentation which are proximity to the actual topological features of the foreground objects. A pioneer work [59] uses an ImageNet-pretrained VGG [60, 27] network after the segmentation network. The perceptual features extracted from different layers of the pretrained VGG network is shown to be approximations of the topological features. The topological loss is then computed between the ground truth and prediction perceptual features. Another topological segmentation loss function [61] was proposed exclusively for the the segmentation of glands. The loss is computed according to the iterations needed to derive the gland skeletons by the image erosion operation. [62] proposed to approximate the topological features of tubular structures by a relationship between the foregrounds in the ground truth and prediction and their soft-skeletons. Using such a relationship as a loss function is shown to strengthen the connectivity along the tubular structure thus providing better topological accuracy.
@@
Another type of methods use strict topological features of the segmentation which are based on the computational topology theory and receieved more research attentions in recent years. The topological features are mainly based on the persistent homology theory [63, 64, 65, 66, 37, 67, 68, 69]. A revolutionary work that first introduce this concept in image segmenation is [63], which computes the persistent features from the ground truth and prediction, matching the features using a Wasserstein distance [70] and use the distance as the loss function. The following works share similar idea. [64] also uses the Wasserstein distance but apply the method on 3D point cloud data. [67] proposed to match the longest persistent barcodes in the prediction to the ground truth. And [37] proposed to use a Hausdorff distance [71] to matches the persistent diagrams to improve the sensitivity of the matching to the outliers. [69] and [68] both aim to compute more precise persistent feature matching, as in earlier methods the actual matchings results are highly inaccurate due to only considering the global topological features. [69] and [68] instead allow the loss to act more selectively on the topological features thus improve the efficiency of the topological loss function. [68] proposed to use an Induced Matching [72] method which constructs a common ambient space between the ground truth and prediction and maps the persistent features in between to find a better matching. However, the computational cost of this method is very high, especially for large images. And the retinal images usually have high resolutions. Alternatively, [69] proposed to leverage the spatial locations of the persistent features referencing their critical cell locations for more faithful persistent feature matching and is computationally much more efficient than [68].
@@ III-B Image Translation Network
Multiple GAN-based or Diffusion model-based image translation methods are explored for our task. For Diffusion-based method, we experiment with StyleDiffusion [58] and Unpaired Neuron Schrodinger Bridge (UNSB) [53]. For GAN-based method, we experiment with CycleGAN [41]. We will show in section IV-D2 that CycleGAN, despite being a classic method, outperforms the other two most recent methods in most datasets. Therefore, we pick CycleGAN as the backbone of our image translation model. Consequently, the following sections will be discussing the implementation based on CycleGAN and we refer the readers to [58, 53] for details about the other two methods and to section IV-D2 for our implementation details.
@@
For the opposite image batch $I_{B}$, the generators (B+A) and discriminator (B) are updated in the same way, as illustrated in Fig. 3. To be more specific, we first update the generators for both $I_{A}$ and $I_{B}$, and then update the discriminators.
@@ III-C2 Persistent Topological Loss
where $s_{p}=\left\|c_{b}(p)-c_{b}(\eta(p))\right\|^{q}$ is a spatial weight that is computed between the creation cells of the persistent features. $p$ and $\eta(p)$ are a persistent feature from $I_{pred}$ and a candidate persistent feature from $I_{GT}$ to be matched with. $q$ is the norm and in practice we take $q=2$. After finding the optimal matching by solving the optimal transport problem, the persistent topological loss function is computed as (reformulating $p$ and $\eta(p)$ by their creation and destruction values):

### v2

@@ Abstract
We identify two major limitations in the existing studies on retinal vessel segmentation: (1) Most existing works are restricted to one modality, i.e, the Color Fundus (CF). However, multi-modality retinal images are used every day in the study of retina and retinal diseases, and the study of vessel segmentation on the other modalities is scarce; (2) Even though a small amount of works extended their experiments to limited new modalities such as the Multi-Color Scanning Laser Ophthalmoscopy (MC), these works still require finetuning a separate model for the new modality. The finetuning will require extra training data, which is difficult to acquire. In this work, we present a foundational universal vessel segmentation model (UVSM) for multi-modality retinal images. Not only do we perform the study on a much wider range of modalities, but we also propose a universal model to segment the vessels in all these commonly-used modalities. Despite being much more versatile comparing with existing methods, our universal model still demonstrates comparable performance with the state-of-the-art finetuned methods. To the best of our knowledge, this is the first work that achieves cross-modality retinal vessel segmentation and also the first work to study retinal vessel segmentation in some novel modalities.  [^1] Our codes and model will be available at the time of publication. We are also actively looking for approaches to release the three new vessel segmentation datasets. However, since retinal images and retinal vessels contain sensitive personal information, we can not make promise at this moment.
@@
retina, retinal images, vessel segmentation, image translation, topology-preserving image segmentation, foundation model.
@@ I Background and Introduction
Retinal vessel segmentation is a fundamental task in retinal image processing. It has multiple useful applications, including retinal image registration [1, 2], human biometric identification [3], retinal artery-vein classification [4] and tree topology analysis in digital images [5]. Clinically, the evaluation of retinal vessels is also important since the vessel features are predictive in several systemic diseases [6, 7, 8] and vessel segmentation is the key towards the automation of the evaluation. However, most works in retinal vessel segmentation focus on only the Color Fundus (CF) images. CF uses white visible light to illuminate the retina and the reflected light is used for generating the image. Nevertheless, although CF was the most widely-used modality in the past, it requires dilation before the imaging and is uncomfortable for the patients. Additionally, the pathological information provided by CF is limited. Therefore, in more recent years, CF is used less frequently and emerging new imaging techniques become increasingly popular. These modalities include (Fig. 1): (1) Multi-Color Scanning Laser Ophthalmoscopy (MC), which uses more intense laser beam and allow exposing structures in deeper layers of the retina; (2) Florescence Angiography (FA), which uses florescence dye to illuminate the retinal vessels and are commonly used in the diagnosis of diseases related to neovascularization; (3) Fundus AutoFlorescence (FAF), which does not rely on external light source but uses natural fluorescence of the retina and provides insights into the health and metabolism of the retinal pigment epithelium (RPE) and photoreceptor cells; (4) Infrared/Near-Infrared Reflectance (IR/NIR), which uses Infrared or Near-Infrared lights to illuminate the image and allows seeing deeper features on the retina such as RPE abnormalities. (5) Optical Coherence Tomography Angiography (OCTA), which uses low-coherence interferometry to measure the backscattered light from retinal layers. It is primarily used to image the micro-vascular structure on the retina. Nevertheless, OCTA has a much smaller imaging area than the other modalities and has significantly different retinal vessel morphology and topology. Therefore, in this work we focus on the first four modalities and the CF.
@@
Figure 1: Overview of multi-modality retinal images and the segmentation results from our UVSM. In each image, the upper half is the original image and the lower half is the segmentation. In the middle, the camera (left) and modality (right) of the image is labeled. ‘Heidel’ refers to Heidelberg Spectralis.
@@
One major challenge that impedes the development of machine learning-based retinal vessel segmentation algorithms is the lack of data. Due to the complexity of the retinal vessels and the existence of many fine details, labeling is very time-consuming. Existing retinal vessel segmentation datasets (especially considering high label quality) are scarce and small even for the CF modality, which usually only have 20-40 images. In addition, to the best of our knowledge, there are no publicly available vessel segmentation datasets for FAF and IR. For FA, there is only a very small public dataset with 8 images [9]. In this work, we prepare three high-quality datasets for FA, FAF and IR, respectively. Each dataset has 40 images and the vessels are carefully labeled by retina experts. Nevertheless, despite the availability of the new datasets, in this work, we only use them for evaluation purpose. The reason is that: we believe that there can never be enough annotated images to allow training a powerful universal model. Instead, we need to train the model with limited existing data and seek to overcome the challenge through innovations in the algorithm. In fact, our segmentation network is trained with only 19 annotated CF images from the widely-used DRIVE [10] dataset.
@@
There are limited previous works which are extended beyond CF modality [11, 12, 13]. In addition to lacking comprehensive modalities (compared to our approach), one still needs to finetune a separate model on the target modality where each model only works for one image modality. We instead aim to build a single universal model that achieves robust segmentation for all the commonly-used modalities. The model will serve as an important tool in cross-modality retinal image study, which becomes increasingly important in modern research of the retina. To achieve the universal segmentation with limited labeled data, we develop our model in two major steps (Fig. 2). Firstly, we propose to use image translation as a preliminary domain adaptation method and train a model that translate input images from arbitrary modalities and cameras into a uniform Topcon (The Topcon TRC-50DX series is explicitly referred to as the gold standard in color fundus photography) CF image modality. We accessed abundant of de-identified multi-modality retinal images from multiple widely-used cameras (Topcon, Heidelberg Spectralis, Optos) at the Jacobs Retina Center at Shiley Eye Institute, University of California, San Diego (UCSD). Those images allow us to train a useful image translation model in a self-supervised approach. Furthermore, we adopt data augmentation to account for the slight domain difference with the other CF cameras which our center does not have, including Canon and Zeiss. Secondly, we train a segmentation model on the Topcon CF domain to perform the segmentation. However, image translation can not fully eliminate the domain gaps between different modalities given such a difficult many-to-one translation task. To solve this problem, we propose to use a topology-aware segmentation method that learns topological features of the vessels in addition to their conventional pixel-level features. The topology-aware segmentation methods were initially proposed to improve the topological accuracy of the segmentation of tubular structures within a certain type of image. We extend these methods to our task and found that they improve the topological accuracy in the segmentation (e.g., preserving vessel continuity, branching and loop features), as well as they force the segmentation to be topologically correct so when images with slight domain gap are inputted, the model can still make the correct segmentation.
@@
Our universal vessel segmentation model is evaluated on 7 diverse datasets (including 3 new datasets we prepared for this work) of the 5 most commonly-used retinal image modalities (CF, FA, FAF, MC, IR/NIR) taken from 5 most commonly-used cameras (Topcon, Canon, Zeiss, Heidelberg Spectralis, Optos). Our universal model is compared with state-of-the-art methods which are finetuned on each of the 7 datasets, respectively, and is shown to achieve comparable performance. Extensive ablation studies were also performed to investigate how different design choices in the proposed pipeline affects the performance of the universal vessel segmentation. The main contributions of this paper are:
@@ II-A Retinal Vessel Segmentation
Retinal vessel segmentation classifies each pixel in the retinal images into foreground (vessel) and background (non-vessel). Current works focus on segmenting vessels in the CF images, where various machine learning-based or non-machine learning-based methods were proposed. A majority of the early methods are based on the matched filter [14, 15, 16], which convolve retinal images with predefined or learned kernels and then adaptively threshold the image to obtain the vessel segmentation map. Another type of methods uses tracking algorithms to mathematically estimate the growth of the retinal vessel trees from seed points [17, 18]. The segmentation can then be derived by extending around the estimated vessel centerlines using the estimated vessel diameters. Some methods also adopt morphological image processing methods such as the top-hat operation to extract the vessel or as a post-processing step [19, 20]. Furthermore, numerous traditional machine learning-based methods were also applied in retinal vessel segmentation such as the support vector machine [21], random forest [22] and Adaboost [23] for supervised methods, and Gaussian mixture model [24], Fuzzy C-Means [25] for unsupervised methods. In more recent years, deep learning-based methods dominate the literature. They mainly rely on the Convolutional Neural Networks (CNN) as a backbone model due to the existence of vast fine details and the spread-out characteristic of the vessels. An introductory work is the Deep Retinal Image Understanding (DRIU) [26], which adopts a VGG [27] network structure as encoder and assigns specialized output layers to the embedded features to segment the vessels and the optic disc from the retina. Based on a widely-used U-Net [28] network structure and to seek further improvement for retinal vessel segmentation, [29, 30, 31, 32, 33] proposed to add various global reasoning modules at the bottom of the network. [32, 33, 34, 35, 31] proposed to add feature aggregation modules in the residual connection at each level of the network. [36, 37] proposed to use deformable convolution where the convolution kernel shapes are learned to adapt to the elongated morphology of the retinal vessels. With similar motivation, [38] proposed an orientation-selective convolution to learn selective receptive fields of the kernel to adapt to the shape of the vessels. [31] proposed to use improved feature fusion methods to better preserve the fine details of vessels in the segmentation. In addition to a single U-Net-like segmentation network, several works [39, 38, 40] proposed to use cascaded networks or a single looped network to iteratively refine the segmentation.
@@ II-C Topology-Aware Image Segmentation
Topology-aware image segmentation aims to preserve the topological accuracy in the segmentation of tubular and net-like structures. The literature mainly considers a topological loss function that extracts topological features from the ground truth and segmentation and minimizes the difference between them. One type of methods extracts certain features from the segmentation which are proximity to the actual topological features of the foreground objects. A pioneer work [59] uses an ImageNet-pretrained VGG [60, 27] network after the segmentation network. The perceptual features extracted from different layers of the pretrained VGG network is shown to be approximations of the topological features. The topological loss is then computed between the ground truth and prediction perceptual features. Another topological segmentation loss function [61] was proposed exclusively for the segmentation of glands. The loss is computed according to the iterations needed to derive the gland skeletons by the image erosion operation. [62] proposed to approximate the topological features of tubular structures by a relationship between the foregrounds in the ground truth and prediction and their soft-skeletons. Using such a relationship as a loss function is shown to strengthen the connectivity along the tubular structure thus providing better topological accuracy.
@@
Another type of methods uses strict topological features of the segmentation which are based on the computational topology theory and received more research attentions in recent years. The topological features are mainly based on the persistent homology theory [63, 64, 65, 66, 37, 67, 68, 69]. A revolutionary work that first introduce this concept in image segmenation is [63], which computes the persistent features from the ground truth and prediction, matching the features using a Wasserstein distance [70] and using the distance as the loss function. The following works share similar idea. [64] also uses the Wasserstein distance but apply the method on 3D point cloud data. [67] proposed to match the longest persistent barcodes in the prediction to the ground truth. [37] proposed to use a Hausdorff distance [71] to match the persistent diagrams to improve the sensitivity of the matching to the outliers. [69] and [68] both aim to compute more precise persistent feature matching, as in earlier methods the actual matching results are highly inaccurate due to only considering the global topological features. [69] and [68] instead allow the loss to act more selectively on the topological features thus improve the efficiency of the topological loss function. [68] proposed to use an Induced Matching [72] method which constructs a common ambient space between the ground truth and prediction and maps the persistent features in between to find a better matching. However, the computational cost of this method is very high, especially for large images, where typical retinal images usually have high resolutions. Alternatively, [69] proposed to leverage the spatial locations of the persistent features referencing their critical cell locations for more faithful persistent feature matching and is computationally much more efficient than [68].
@@ III-B Image Translation Network
Multiple GAN-based or Diffusion model-based image translation methods are explored for our task. For Diffusion-based method, we experiment with StyleDiffusion [58] and Unpaired Neuron Schrodinger Bridge (UNSB) [53]. For GAN-based method, we experiment with CycleGAN [41]. We will show in section IV-D2 that CycleGAN, despite being a classic method, outperforms the other two most recent methods in most datasets. Therefore, we pick CycleGAN as the backbone of our image translation model. Consequently, the following sections discuss the implementation based on CycleGAN and we refer the readers to [58, 53] for details about the other two methods and to section IV-D2 for our implementation details.
@@
For the opposite image batch $I_{B}$, the generators (B+A) and discriminator (B) are updated in the same way, as illustrated in Fig. 3. In addition, we first update the generators for both $I_{A}$ and $I_{B}$, and then update the discriminators.
@@ III-C2 Persistent Topological Loss
where $s_{p}=\left\|c_{b}(p)-c_{b}(\eta(p))\right\|^{q}$ is a spatial weight that is computed between the creation cells ($c_{b}(\cdot)$) of the persistent features. $p$ and $\eta(p)$ are a persistent feature from $I_{pred}$ and a candidate persistent feature from $I_{GT}$ to be matched with. $q$ is the norm and in practice we take $q=2$. After finding the optimal matching by solving the optimal transport problem, the persistent topological loss function is computed as (reformulating $p$ and $\eta(p)$ by their creation and destruction values):